# Insight-HXMT View of the BHC Swift J1727.8-1613 during its outburst in 2023

### v2

@@ 4.2 Evolution of the physical configuration and spectral nature
Looking at the spectral parameters, and also the location of the shock, we designate that the source was already in an intermediate state at the start of our analysis period. As the outburst progressed, it was transitioning towards softer states. From the values of the parameter $\cos(i)$ of the pexrav model, we find that the inclination of the disk was $>60^{\circ}$, which makes this source a highly inclined system. The hydrogen column density ($N_{H}$) showed a variation, which is common for Galactic black hole binaries unless the system is heavily obscured, which has recently been found for the newly discovered BH Swift J151857.0-572147 (Mondal et al. 2024). The average $N_{H}$ has an average value of $0.27_{-0.17}^{+0.08}\times 10^{22}$ cm-2.
@@ 5 Summary and Conclusions
(i) Quasi-periodic oscillations were present during the entire period of our analysis, which showed evolution in QPO frequency.
(ii) The $Q$-factor and $RMS$ help designate the nature of the QPOs as type-C.
(iii) QPOs originated from the shock instabilities that was formed by matter having an angular momentum distribution deviated from the Keplerian distribution. As the shock moved inwards, the QPO frequency was observed to increase.
@@
(vii) The hydrogen column density varied in the range of $0.10-0.35\times 10^{22}$ cm-2, which is quite common for interstellar absorption of Galactic black holes.
@@ 7 Acknowledgements
KC acknowledges support from the SWIFAR postdoctoral fund of Yunnan University. SM acknowledges the Ramanujan Fellowship (# RJF/2020/000113) by SERB-DST, Govt. of India for this research. CBS is supported by the National Natural Science Foundation of China under grant no. 12073021.

### v3

@@ 4 Discussion
The photon index of power-law ($\Gamma$) is an important property to decide the spectral nature of an ongoing outburst. This decides the slope in the spectrum, which tells whether the spectrum extends up to high energy or not. When an outburst stays in the hard state, the photon index of power-law ($\Gamma$) generally stays low $\sim 1.7$ (Remillard & McClintock 2006). From the expression of power-law component ($A(E)\propto E^{-\Gamma}$), this suggests that the spectrum extends up to high energy. As the source gets towards intermediate states, $\Gamma$ becomes $\sim 2$ (e.g., Remillard & McClintock 2006), even up to 3, sometimes. In the high soft state, $\Gamma$ reaches a value of $>3$ (Remillard & McClintock (2006), for a review), although sources have been discovered to reach soft state with $\Gamma<3$. According to the TCAF configuration (Chakrabarti & Titarchuk 1995), as discussed earlier in the Introduction, the shock location, that produces both the change in the spectral and timing properties, changes in size in different spectral states. In the HS it stays very far from the source, at $\sim 1000~{}r_{s}$. The disk is even further away. As a result, only a few photons come in the process of radiation. As the disk approaches closer, the $X_{s}$ becomes smaller and the component from the disk starts to gradually dominate over the hard power-law component. In the intermediate states, these two components stay comparable. This time the shock stays at a few $100~{}r_{s}$. In the SS, the disk moves even closer and $X_{S}$ becomes very small (a few tens of $r_{s}$). Thus, the disk component now fully dominates, and the spectrum becomes very soft. Thereby, the location of the shock determines the spectral state of the outburst. Looking at the spectral parameters, and also the location of the shock, we propose that the source was already in an intermediate state at the start of our analysis period. As the outburst progressed, it was transitioning towards softer states. From the values of the parameter $\cos(i)$ of the pexrav model, we find that the inclination of the disk was $>60^{\circ}$, which makes this source a highly inclined system. This is in agreement with the recent reports of Bouchet et al. (2024), Yu et al. (2024). The hydrogen column density ($N_{H}$) showed a variation, which is common for Galactic black hole binaries (Eckersall et al. 2017; Chatterjee et al. 2023) unless the system is heavily obscured, which has recently been found for the newly discovered BH Swift J151857.0-572147 (Mondal et al. 2024), Swift J1658.2-4242 (Mondal & Jithesh 2023 and references therein). The $N_{H}$ has a variation in the range of $(0.12\pm 0.02-0.39\pm 0.08)\times 10^{22}$ cm-2.
The QPO frequency can be produced from the oscillation of the Compton corona, which can also be the origin of the hard power-law photons that extend the spectrum to higher energies (e.g., Lee & Miller 1998; Kumar & Mishra 2014; Karpouzas et al. 2020; Liu et al. 2022; Bellavita et al. 2022; etc.). Therefore, these two features might be interlinked or correlated. As described in the introduction, the oscillation of the CENBOL can be the mechanism producing the QPOs. This CENBOL is also responsible for producing the power-law part of the spectrum. The change in the size of the shock location gives the change of spectral states as well as the frequency of the QPOs. We found such correlations for the source Swift J1727.8-1613 during its 2023 outburst.
@@ 5 Summary and Conclusions
(i) Quasi-periodic oscillations were present during the entire period of our analysis, which showed evolution in QPO frequency from $0.21\pm 0.01$ to $1.86\pm 0.01$ Hz within this analysis period.
(ii) The values of the $Q$-factor and $RMS$ best represent the QPOs as type-C.
(iii) Our study infers that the QPOs might have originated due to the resonance oscillations of the shocks. As the shock moved inwards (outwards), the QPO frequency increased (decreased).
@@
(vii) The hydrogen column density varied in the range of $(0.12\pm 0.02-0.39\pm 0.08)\times 10^{22}$ cm-2., which is quite common for interstellar absorption of Galactic black holes.
@@ 7 Acknowledgements
We thank the scientific editor and the anonymous referee(s) for their detailed comments and insightful suggestions that improved the quality of the paper. KC acknowledges support from the SWIFAR postdoctoral fund of Yunnan University. SM acknowledges the Ramanujan Fellowship (# RJF/2020/000113) by SERB/ANRF-DST, Govt. of India for this research. CBS is supported by the National Natural Science Foundation of China under grant no. 12073021.